OpenStudy (anonymous):

A gas cylinder contains only the gases radon, nitrogen, and helium. The radon has a pressure of 222 torr while the nitrogen has a pressure of 446 torr. If the total pressure inside the cylinder is 771 torr, what is the pressure that is due to the helium?

OpenStudy (anonymous):

P(cylinder) = P1 + P2 + P3 (gases) so 771 - 222 - 446 = P(He) P(He) = 109.. assuming ideality of gases...

OpenStudy (jfraser):

dalton's law of partial pressures: the total pressure of a system of gases in a container is equal to the partial pressures of each component gas. You know three of the four gases, and the total pressure.